OpenStudy (anonymous):

1/81 = b^(-4) I know the answer is 3, but I need to show the work

OpenStudy (anonymous):

help please!!!!!!

OpenStudy (anonymous):

will medal (if i can figure out how)

OpenStudy (anonymous):

\[b^{-4} = \frac{ 1 }{ b^4 }\] \[\frac{ 1 }{ 81 } =\frac{ 1 }{ b^4 }\] \[81 = {b^4}\] \[\sqrt{81} = \sqrt{b^2}\] \[9 = b^2\] \[\sqrt{9} = \sqrt{b}\] b = 3

OpenStudy (anonymous):

thank you thank you thank you!

OpenStudy (anonymous):

anytime :)

OpenStudy (anonymous):

@sangya21 okay, real quick, how did u get b^2 out of b^4 on line four?

OpenStudy (anonymous):

\[81 = b^2 *b^2\] \[\sqrt{81} =\sqrt{b^4}\] \[9 ={b^2}\]

OpenStudy (anonymous):

@sangy21 i see it now. u worked down the roots

OpenStudy (anonymous):

Yups.